1. A sandwich shop owner has the following information: P = MR = $ 4, ATC = $ 2, AVC
= $ 1, MC = 4, and Q = 500. From this, she can determine:
a. her profits are not being maximized.
b. she has earned economic profits of $ 1,000.
c. she has earned economic profits of $ 1,500.
d. she should sell fewer sandwiches.

2. If a firm shuts down in the short run, it will:
a. incur losses equal to its fixed costs.
b. produce at the output level where MC = MR.
c. reduce its losses to zero.
d. do this because P > AVC.
1
Expert's answer
2015-06-30T00:00:39-0400
1. A sandwich shop owner has the following information: P = MR = $ 4, ATC = $ 2, AVC = $ 1, MC = 4, and Q = 500. From this, she can determine:
As MR = MC and P>ATC
TP = (P - ATC)*Q = (4 - 2)*500 = $1000, so:
b. she has earned economic profits of $ 1,000.
2. If a firm shuts down in the short run, it will:
a. incur losses equal to its fixed costs.